tom has 42 coins in nickels and quarters. He has 6 more quarters than nickels, if x= number of nickels and y= number of quarters, then what is the solution set for the number of nickels and quarters?
Answer by stanbon(75887)

tom has 42 coins in nickels and quarters. He has 6 more quarters than nickels, if n = number of nickels and q = number of quarters, then what is the solution set for the number of nickels and quarters?
---------------
n + q = 42 coins
q = n+6
------
Substitute for "q" and solve for "n"::
n + n+6 = 42
2n = 36
n = 18 (# of nickels)
-----
q = n+6 = 18+6 = 24 (# of quarters)
--------------
Cheers,
Stan H.
-------------